15= (2*0.45)+(x*0.16)+3.00
Answer by stanbon(75887)

15= (2*0.45)+(x*0.16)+3.00
------
15 = 0.90+0.16x+3
------
12 = 0.9 + 0.16x
------
11.1 = 0.16x
-----
x = 11.1/0.16
x = 69.375
-----
Cheers,
Stan H.
-------------